OpenStudy (anonymous):

An expression is shown below: square root of 50 plus square root of 2 Which statement is true about the expression?

OpenStudy (anonymous):

It is irrational and equal to 2 multiplied by the square root of 13. It is irrational and equal to 6 multiplied by the square root of 2. It is rational and equal to 2. It is rational and equal to 6.

OpenStudy (gorv):

we know 2 is not rational so it will be irrational

OpenStudy (anonymous):

I think its a

OpenStudy (gorv):

\[\sqrt{50}+\sqrt{2}=\sqrt{25*2}+\sqrt{2}=5*\sqrt{2}+\sqrt{2}=6*\sqrt{2}\]

OpenStudy (igreen):

I'm going with B. They're both irrational and it equals \(6 \cdot \sqrt 2\)

OpenStudy (gorv):

@briannamooney now u tell me the ans ??

OpenStudy (anonymous):

It is possible to simplify 50 as represented in a square root. Factoring 50, we get: 2 * 5 * 5

OpenStudy (anonymous):

\[\sqrt[6]{2}\]

OpenStudy (anonymous):

We've factored 50 as far as it will go. You might notice that there are two fives in 50. What happens if we multiply those two fives together? We get 25 as a result of that and 25 so happens to be a perfect square. You can take perfect squares out of the square root symbol. Similarly, find for 2 and add it up..!

OpenStudy (anonymous):

I understood what he did until he multiplied \[\sqrt[5]{2}\] and got 6 , i still need to look over how to solve a radical